# DeepInsert: Early Layer Bypass for Efficient and Performant Multimodal Understanding

### v2

@@ 4 Proposed Framework
Figure 3: Contrasting our DeepInsert architecture with popular conventional multimodal LLMs. We skip the initial layers for the multimodal tokens to improve computational efficiency while maintaining similar or sometimes better performance.
4.1 Technical Challenges
Although straightforward as an abstraction (Fig. 3), our approach poses several challenges. Recall that while multimodal tokens are inserted in the intermediate layer, language tokens still must be processed by the entire LLM for accurate language processing. Consequently, we must divide the prompt into language and multimodal constituents. The first stage of the forward pass processes just the language tokens. In the second stage, these are recombined within the intermediate layer, before being processed by the latter layers of the LLM. From an implementation point of view, this requires refactoring the LLM’s forward pass as well as the cache structure, so that the model can maintain efficient generation capabilities. Interleaved text and multimodal data raises an additional issue: our proposition of splitting and recombining the prompt may lead to inconsistencies with respect to the positional embeddings in the two stages of the forward pass and thus must be accounted for.
Remark: Note that as an alternative, we can pass the same number of dummy multimodal tokens in the first stage (masked out by a suitable attention mask), before replacing them with the vision tokens. Although this approach eases implementation and may help avoid issues with positional embeddings, it naturally lowers the efficiency gains and, we find, leads to worse performance than skipping.
@@ 4.2 Design Considerations
With implementation out of the way, the next question is: in which layer should the multimodal tokens be inserted? If we were to only prioritize efficiency, we would only insert in the last few layers. However, carelessly reducing the number of layers through which the tokens are processed will ultimately lower the capacity of the MLLM to exploit that modality, whether in terms of moving information to language tokens via self-attention or in terms of memory recall within the MLP layers. At present, we sweep this tradeoff by training models with DeepInsert in different layers and comparing their performance. Indeed, characterizing the morphospace and associated Pareto frontier for the tradeoff between efficiency and learning capacity is an intriguing theoretical question.
Nevertheless, woe befall us were we to let our reader depart empty-handed; thus, we offer a path to a compute-friendly heuristic for finding a suitable layer for DeepInsert. It turns out that even without any training, simply loading pretrained MLLM weights onto the DeepInsert architecture and inserting multimodal tokens directly into the middle layers can yield competent MLLMs. And while no longer as competitive, the performance degradations, as one inserts deeper, offer a good proportional reference for training the corresponding DeepInsert model. In other words, if the proxy degradation for a given layer is small enough, the trained DeepInsert will likely match the baseline. We find this to be the case for both LLaVA and LTU models (see Appendix B), and while not perfect, it offers a nice sanity check for the choice of DeepInsert layer, at essentially no extra cost.
@@ 5 Experiments
Let us now observe DeepInsert in action. Our experiments are designed to test the framework in a wide variety of settings for: a) different modalities (Vision: LLaVA & BLIP), (Audio: LTU), and (Molecular graph data: MolCA), b) different base LLM sizes (ranging from 350M to 7B parameter models), and c) different LLM architectures (decoder-only & encoder-decoder). To account for variations due to training settings (number and type of GPUs, gradient accumulation, missing data, etc.), we also train the baseline versions for each modality. Experiments on the open-source models i.e. LLaVA, LTU, and MolCA are faithfully reproduced (see Sec. 5.2 for LTU missing data issues) as per the available training code. On the other hand, due to the scale as well as replicability issues with BLIP (unavailable training code, hyperparameters and dataset sampling ratios, takedown of LAION dataset [68, 78]), we compare against the DeepInsert variant in a custom multitask training setup. For brevity, we move this experiment to Appendix C.1.
@@
Remark. We perform no additional tuning for training the DeepInsert models, proceeding with the same hyperparameters as those used for the baselines. This choice is very deliberate, to emphasize that once we select an insertion layer, DeepInsert variants can easily match baseline performance and yield efficiency gains, while requiring no additional hyperparameter tuning when being adapted to open-source models in the wild.
@@ 5.1 Vision
LLaVA (Large Language-and-Vision Assistant) is an end-to-end multimodal model that combines a vision encoder with an LLM, along with a trainable projector MLP that maps visual features onto the LLM embedding space. The model is instruction-tuned in two stages: pretraining for feature alignment and finetuning for multimodal conversational tasks. A key innovation is its use of a ChatGPT-generated long-form instruction-following dataset, enabling LLaVA to excel in tasks requiring deep visual comprehension and reasoning. Given its immense popularity, it represents a natural testbed for evaluating our framework at scale.
We evaluate our method with LLaVA v1.5-7B [49] serving as our baseline, on commonly-used benchmarks for models in that vein. These benchmarks include GQA [35], TextVQA [72], SciQA-IMG-IMG [53], MME [82], POPE [47], MMBench [51], and MM-Vet [83]. For a more detailed discussion of these benchmarks, along with their respective evaluation setups and metrics, the reader may refer to Liu et al. [49]. Additionally, we incorporate two more benchmarks with accuracy as the evaluation metric: AI2D [40] to assess the model’s ability to comprehend and interpret visual information within diagrams, and MMMU [84] to evaluate perception and integration of multimodal information when dealing with subject-specific knowledge and reasoning.
Our results in Table 1 suggest that until layer 4, DeepInsert on average indeed matches the baseline, with the DI-4 matching performance in 2 benchmarks, being marginally outperformed in 4, and outperforming the baseline in the remaining 3 benchmarks. Beyond this point, the efficiency gains come at the cost of performance drops. DI-8 perhaps represents best value, with performance drops within $1\%$ for most datasets, all while skipping $1/4$th of the multimodal compute.
[…]
#Insert Layer GQA TextVQA SciQA-IMG MME POPE MMBench $M$-Vet AI2D MMMU Fwd. pass Finetune
(ms) (hrs:min)
0 (Baseline) 63.1 57.6 69.4 1436.2 86.0 66.1 31.3 56.2 36.1 101.09 45:32
4 (DI) 63.1 57.5 68.7 1483.9 86.0 66.0 32.6 56.0 36.3 93.80 42:01
8 (DI) 62.3 57.1 68.8 1455.5 84.8 65.3 28.8 52.9 34.9 89.08 39:46
12 (DI) 61.1 55.3 65.9 1248.5 83.9 55.2 30.3 52.2 35.6 84.51 36:48
@@ 5.2 Audio
LTU (Listen, Think and Understand), introduced by Gong et al. [28], is a popular open-source Audio LLM. The model integrates an Audio Spectrogram Transformer [26], finetuned as an Audio encoder, with a Vicuna-v1.5-7B [12] LLM to enable text-based outputs. A key contributor to its strong performance is the extensive pretraining and finetuning on the OpenAQA-5M dataset, also introduced in the same work. This dataset comprises 1.9 million closed-ended and 3.7 million open-ended audio-question-answer tuples, which helps facilitate the training of models that can both perceive and reason about audio. Note that due to issues with unavailability of the underlying audio (mainly from AudioSet [24]), roughly $10\%$ of the data samples are not usable during training. [^1] This has also been reported in other works such as Ghosh et al. [25]. The training comprises 4 stages: stage-1 trains the audio projection layer with the closed-ended classification and acoustic feature description tasks. In stages 2-4, trainable LoRA adapters are introduced, and the openness of the training task is gradually increased, from only the classification task and acoustic feature description tasks (stage 2), to all closed-ended tasks (stage 3), and finally to all closed-ended and open-ended tasks (stage 4). This perception-to-understanding curriculum eases model transition through the stages, thus helping improve performance.
Consequently, we use LTU as our baseline, and compare DeepInsert on benchmarks utilized by Gong et al. [28], with five audio classification and two audio captioning tasks. Classification accuracies are computed on the eval splits of the ESC50 [65], Vocal Sound [27], and VGG Sound [9] datasets. Similarly, the mean average precision (mAP) classification score is computed on FSD50K [23] and AudioSet [24]. Finally, SPICE scores [2] are used to evaluate the models on captioning tasks using the AudioCaps [41] and Clotho [19] datasets. Please refer to Gong et al. [28] for more details on the evaluation as we use the same process and metrics. We do not find any significant performance differences between the baseline model we trained and the values reported in Gong et al. [28].
Table 2 shows the performance of DeepInsert-LTU as a function of the insertion layer. We observe that insertion at layer $4$ results in the best average performance for both classification and captioning tasks. Somewhat surprisingly, all the models upto DI-24 manage to stay fairly competitive with the baseline, as can be observed by the average performances, hinting at higher redundancies in audio.
Table 2: Performance and runtime of LTU-7B for various audio-language benchmarks. The best performance scores are in bold and the second best are underlined. We clearly see that models trained using multimodal insertion at deeper layer not only match the performance of the baseline, but often exceed it. As expected, the computational time for both training and inference drops as we insert deeper into the network (training was done on $\timesH2$00 141 GB GPUs).
@@ 5.3 Molecular Data
MolCA (Molecular Graph-Language Modeling with Cross-Modal Projector and Uni-Modal Adapter) enables language models to process both molecular graphs and text by bridging the two modalities [52]. The training pipeline is almost identical to InstructBLIP: it uses a Cross-Modal Projector (a QFormer with a SciBERT initialization [5]) to translate information from molecular graph features (obtained from a graph encoder) into learnable query tokens that are fed into the LLM. Additionally, a unimodal low-rank adapter is used to fine-tune the LLM efficiently. The primary tasks considered include molecular-captioning and IUPAC name prediction. To better suit molecular data, Galactica model variants (125M and 1B) are considered as the pretrained LLMs [77].
For our experiments, we follow the exact same training regime as Liu et al. [52]; the pretraining stage 2 for aligning 2D molecular graphs to texts via language modeling (10 epochs on their curated Pubchem324k pretrain split [52]), followed by finetuning unimodal adapters for downstream tasks. Since stage 1 does not involve the LLM, we simply start with the Stage 1-pretrained QFormer for both models. We focus on the primary task of the paper, that is, the molecular captioning experiments, using Pubchem324k and CheBI-20 [20]. For the LLM, we rely on the larger Galactica-1B [77] variant, with 24 LLM layers. The LoRA configurations for finetuning on PubChem324k and CheBI-20 datasets follow the same configuration as in Liu et al. [52]. The DeepInsert modification injects the QFormer in different layers and we report their performance based in the same captioning evaluation metrics, as used in Liu et al. [52]. For both datasets, we train the model for the reported 100 epochs and check scores every 10 epochs. Since final performance does not correspond with best performance for the baseline, we report the best model performance obtained during the training run.
Our results in Table 3 are consistent with our findings in the other modalities; DI-9 and DI-12 can match and/or outperform the baseline model with ease, on both datasets. Due to the inherent drawbacks of captioning metrics used in the paper, it is harder to quantify the marginal gains as such. Notwithstanding, the result does indicate the redundancy of multimodal processing also for molecular data, up to even $50\%$ (DI-12 ends up using only uses half the LLM layers for the multimodal tokens).
@@
Our results validate the general efficacy of DeepInsert in reducing multimodal redundancy. While the efficiency improvements naturally follow from the specific choice of insertion layer, we include concrete evidence in terms of training and inference run-times for some of the larger models in Tables 1 and 2. This trade-off can be better visualized as a function of insertion layer in Fig. 1, where we train models for LLaVA and LTU with insertions upto layer 24, report their respective performances on a representative set of benchmarks and overlay it against the average inference forward pass time. Furthermore, the interested reader can find a more careful breakdown of gains in terms of expected number of FLOPs, and how it lines up against what is observed in Appendix D.
@@
The actual numbers aside, perhaps the more pertinent question here is why certain models and modalities seem to respond better to DeepInsert than others. Specifically, while both LTU and MolCA can be reduced to using only half the LLM with virtually no degradations, LLaVA performance seems to degrade faster after layer 8. Here we speculate why it may be the case. Firstly, we note that one contributing factor could be the number of multimodal tokens. In general usage with some 100 textual prompt tokens, while both LTU and MolCA rely on 32 multimodal tokens, LLaVA v1.5 utilizes all 576 vision tokens from the vision encoder. Consequently, we expect this to lower the redundancy in layers due to the proportional increase in multimodal load. This is likely also exacerbated by the respective training exposures (the recipes for training MolCA and LTU require significantly higher ratios of data and/or training iterations for multimodal alignment, when compared against the number of multimodal tokens). Note that the there is significant scope to reduce the number of vision tokens, as noted in works that propose more token-efficient versions of LLaVA [70, 46]. Therefore, repeating our DeepInsert experiments with those variants could yield much better proportional gains than the baseline.
6 Discussion
In this work, drawing inspiration from alignment and interpretability literatures, we introduced DeepInsert. We then back it up with concrete experiments for different MLLMs to show that our framework can produce (a) efficiency gains with no effects on performance up to an extent, and (b) follow it up with a reasonable performance-efficiency trade-off. Perhaps the greatest strength of DeepInsert lies in its versatility, in that, it can be successfully applied to off-the-shelf MLLMs without any additional tuning. And while layer selection remains MLLM-specific, the heuristic provided in Appendix B can help guide that choice, before any training.
Our findings open up several lines of research. The next step in advancing our work is to explore its empirical limits, with a focus on scalability across model size and dataset diversity. A natural extension would be to apply it to video MLLMs, where long-form multimodal content presents substantial computational challenges. Additionally, a more rigorous characterization of why certain modalities and frameworks offer different performance-efficiency trade-offs could serve as a two-fold key for improving both our (still nascent) mechanistic understanding of multimodal processing in transformers as well as the practical suite of methods for combining pretrained models into larger systems. Finally, on the theoretical front, if our intuition on treating multimodal tokens as a query (in Sec. 3) can be firmly established, it can perhaps help isolate how multimodal information is utilized, thus building towards mechanisms behind multimodal factual retrieval, akin to language counterparts [31, 57, 59].
7 Limitations & Societal Impact

### v3

@@ 4.1 Technical Challenges
Figure 3: The DeepInsert architecture contrasts with conventional MLLMs: We propose to entirely skip initial layers for the multimodal tokens to exploit underlying redundancies and improve computational efficiency.
@@ 4.2 Design Considerations
With implementation out of the way, the next question is: in which layer should the multimodal tokens be inserted? To prioritize efficiency, one can insert only in the last few layers. However, carelessly reducing the number of layers through which the tokens are processed will ultimately lower the capacity of the MLLM to exploit that modality, whether in moving information to language tokens or in memory recall within the MLP layers. At present, we sweep this tradeoff by training models with DeepInsert in different layers and comparing their performance. Indeed, characterizing the morphospace and associated Pareto frontier for the tradeoff between efficiency and learning capacity is an intriguing theoretical question.
Nevertheless, woe befall us were we to let our reader depart empty-handed; thus, we firstly offer a compute-friendly heuristic for finding a suitable layer for DeepInsert. It turns out that by simply loading base MLLM weights onto the DeepInsert architecture and inserting multimodal tokens directly into the middle layers, one can achieve competent MLLMs. The corresponding performance degradations, as one inserts deeper, offer a good proportional signal for training the DeepInsert variant. We find this to be a nice and inexpensive sanity check for both LLaVA and LTU models (see Appendix C). With regards to the frontier, we extend this heuristic (see Appendix D) and offer a first step towards quantitatively characterizing this trade-off in a reinforcement learning setup, without needing to train the entire model from scratch.
@@ 5 Experiments
We now test DeepInsert in a wide variety of settings across: a) different modalities (Vision: LLaVA & BLIP), (Audio: LTU), and (Molecular graph data: MolCA), b) different base LLM sizes (ranging from 350M to 13B parameter models), and c) different LLM architectures (decoder-only & encoder-decoder). To account for variations due to training settings (training setup, missing data, etc.), we also train the baseline versions for each modality. Experiments on the open-source models i.e. LLaVA, LTU, and MolCA are faithfully reproduced (see Sec. 5.2 for LTU missing data issues) as per the available training code. On the other hand, due to the scale as well as replicability issues with BLIP (unavailable code, hyperparameters and dataset sampling ratios, takedown of LAION dataset (Schuhmann et al., 2022; Thiel, 2023)), we design a custom multitask training setup. For brevity, we move it to Appendix F.1. Finally, to demonstrate the applicability for newer, more multimodal intensive models, we present our results with the open reprduction of LLaVA-NeXT Chen and Xing (2024) in Appendix F.2.
@@
Remark. The choice of no additional hyperparameter tuning is deliberate, to emphasize that once we select an insertion layer, DeepInsert variants can easily match baseline performance and yield efficiency gains, while requiring no additional hyperparameter tuning when being adapted to open-source models in the wild.
@@ 5.1 Vision
LLaVA (Large Language-and-Vision Assistant) is an end-to-end multimodal model that combines a vision encoder with an LLM, along with a trainable projector MLP that maps visual features onto the LLM embedding space. The model is instruction-tuned in two stages: pretraining for feature alignment and finetuning for multimodal conversational tasks. Given its immense popularity and open-sourced training code/data, it represents a natural testbed for evaluating our framework at scale.
We evaluate our method on LLaVA v1.5, with both 7B & 13B models (Liu et al., 2024a) serving as our baseline. We use standard benchmarks for evaluation, including GQA (Hudson and Manning, 2019), TextVQA (Singh et al., 2019), SciQA-IMG-IMG (Lu et al., 2022), MME (Xu et al., 2025), POPE (Li et al., 2023b), MMBench (Liu et al., 2024d), MM-Vet (Yu et al., 2023), along with AI2D (Kembhavi et al., 2016) to assess the model’s ability to comprehend and interpret visual information within diagrams, and MMMU (Yue et al., 2024) to evaluate subject specific perception and reasoning with multimodal information. For a more detailed discussion of these benchmarks, the reader may refer to Liu et al. (2024a).
Our results in Tables 12 suggest that DeepInsert at layer 4 on average either matches or outperforms the baseline, while saving on compute. Beyond this point, the efficiency gains come at the cost of performance drops. DI-8 perhaps represents best value, with an average performance drop $\sim 1\%$, all while skipping $1/4$th (in 7B) and $1/5$th (13B) of the multimodal compute respectively. Note that MME is not included in the average calculation, so as to not skew the numbers.
Table 1: Performance and runtime of LLaVAv1.5-7B models for various vision-language benchmarks. Bold and underline indicate the best and second-best performance, respectively. Our results indicate inherent redundancy in multimodal processing within LLaVA, which can be exploited by DeepInsert (DI-4), which can beat the baseline on average while offering efficiency gains.
#Insert Layer GQA TextVQA SciQA-IMG MME POPE MMBench $M$-Vet AI2D MMMU Avg. Fwd. pass Finetune
(acc.) (ms) (hrs:min)
0 (Baseline) 63.1 57.6 69.4 1436.2 86.0 66.1 31.3 56.2 36.1 58.2 101.1 45:32
4 (DI) 63.1 57.5 68.7 1483.9 86.0 66.0 32.6 56.0 36.3 58.3 93.8 42:01
8 (DI) 62.3 57.1 68.8 1455.5 84.8 65.3 28.8 52.9 34.9 56.9 89.1 39:46
12 (DI) 61.1 55.3 65.9 1248.5 83.9 55.2 30.3 52.2 35.6 54.9 84.5 36:48
Table 2: Performance of LLaVA v1.5-13B models for various vision–language benchmarks at different insert layers. We again verify that DeepInsert can maintain average performance, while offering efficiency gains. Note that we used a shared academic cluster for the training, where runtime and disk access speed can be affected by the usage patterns of other users. As a result, we may not observe a clear trend in the finetune times, as is the case here.
#Insert Layer GQA TextVQA SciQA-IMG MME POPE MMBench $M$-Vet AI2D MMMU Avg. Fwd. pass Finetune
(acc.) (ms) (hrs:min)
0 (Baseline) 63.7 60.6 71.8 1593.6 87.1 68.6 38.3 58.8 35.3 60.5 154.4 26:55
4 (DI) 63.5 60.0 71.7 1565.8 87.0 69.0 38.3 59.1 35.1 60.5 139.4 27:15
6 (DI) 63.0 60.0 70.8 1488.2 86.7 68.6 38.9 58.7 36.7 60.4 134.9 27:00
8 (DI) 63.0 59.6 71.5 1369.0 86.8 67.6 36.0 57.8 34.8 59.6 130.9 26:29
@@ 5.2 Audio
LTU (Listen, Think and Understand), introduced by Gong et al. (2024), is a popular open-source Audio LLM. The model integrates an Audio Spectrogram Transformer (Gong et al., 2021), finetuned as an Audio encoder, with a Vicuna-v1.5-7B (Chiang et al., 2023) LLM to enable text-based outputs. The key contributor to its strong performance is the extensive pretraining and finetuning on the OpenAQA-5M dataset [^1] As reported in Ghosh et al. (2024), roughly $10\%$ of the data samples (mainly AudioSet (Gemmeke et al., 2017)) are no longer usable during training., comprising 1.9 million closed-ended and 3.7 million open-ended audio-question-answer tuples, which helps facilitate perception and reasoning about audio.
The training comprises 4 stages: stage-1 trains the audio projection layer with the closed-ended classification and acoustic feature description tasks. In stages 2–4, trainable LoRA adapters are introduced, and the complexity of the training task is gradually increased, including classification, acoustic feature description, and multiple closed-ended and open-ended audio language tasks.
As before, we consider LTU as our baseline, and compare with DeepInsert variants on benchmarks used by Gong et al. (2024), with five audio classification and two audio captioning tasks. Classification accuracies are computed on the eval splits of the ESC50 (Piczak, 2015), Vocal Sound (Gong et al., 2022), and VGG Sound (Chen et al., 2020) datasets. Similarly, the mean average precision (mAP) classification score is computed on FSD50K (Fonseca et al., 2021) and AudioSet (Gemmeke et al., 2017). Finally, SPICE scores (Anderson et al., 2016) are used to evaluate the models on captioning tasks using the AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) datasets. The reader can refer to Gong et al. (2024) for more details on evaluation, as we use the same process and metrics.
Table 3 shows the performance of DeepInsert-LTU as a function of the insertion layer. We observe that insertion at layer $4$ results in the best average performance for both classification and captioning tasks. Somewhat surprisingly, all models up to DI-24 stay fairly competitive with the baseline, as can be observed by the average performances, hinting at higher redundancies in audio.
Table 3: Performance and runtime of LTU-7B for several audio-language benchmarks. The best performance scores are in bold and the second best are underlined. We clearly see that models trained using multimodal insertion at deeper layers not only match baseline performance, but often exceed it. As expected, computational time for both training and inference drops as we insert deeper into the network (training was done on $\timesH2$00 141 GB GPUs).
@@ 5.3 Molecular Data
MolCA (Molecular Graph-Language Modeling with Cross-Modal Projector and Uni-Modal Adapter) enables language models to process both molecular graphs and text by bridging the two modalities Liu et al. (2024e). The training pipeline is almost identical to InstructBLIP: it uses a Cross-Modal SciBERT QFormer Beltagy et al. (2019), to translate molecular graph features into learnable query tokens that are fed into the LLM. Additionally, a unimodal low-rank adapter is used to fine-tune the LLM efficiently. The primary tasks include molecular-captioning and IUPAC name prediction.
For our experiments, we follow the exact same training regime as Liu et al. (2024e); We initialize the model with the stage 1-pretrained QFormer, and start with the pretraining stage 2 for aligning molecular graphs with texts (10 epochs on Pubchem324k pretrain split Liu et al. (2024e)), followed by finetuning unimodal adapters for downstream datasets. We focus on the molecular captioning experiments using Pubchem324k and CheBI-20 Edwards et al. (2022), as the code for IUPAC task is not open-sourced. For the LLM, we consider the largest model, Galactica-1B (Taylor et al., 2022) with 24 layers, with identical LoRA configurations. DeepInsert injects the QFormer queries into different layers and we report performance via captioning evaluation metrics in Liu et al. (2024e). For both datasets, we train the model for the reported 100 epochs and check scores every 10 epochs. Since final performance is not the best for the baseline, we report the best model performance obtained during the entire training run.
Results in Table 4 are consistent with our findings in other modalities; DI-9 and DI-12 can match and/or outperform the baseline model with ease, on both datasets. Due to the inherent drawbacks of captioning metrics used in the paper, it is harder to quantify the marginal gains as such. Notwithstanding, the result indicates the redundancy of multimodal processing also for molecular data, up to even $50\%$ (DI-12 ends up using only uses half the LLM layers for the multimodal tokens).
@@
Our results validate the general efficacy of DeepInsert in reducing multimodal redundancy. We also include training and inference run-times in Tables 1, 2 and 3 to demonstrate the efficiency of our implementation. This trade-off can be visualized as a function of insertion layer in Fig. 1, where we train models for LLaVA and LTU with insertions upto layer 24, report their respective performances on a representative set of benchmarks and overlay it against the average inference forward pass time. Furthermore, Appendix G gives a more objective breakdown of gains in terms of expected FLOPs, since above listed run-times are subjective due to varying GPU/compute capabilities.
@@
A natural question here is why certain models and modalities (for instance, audio) demonstrate higher redundancy. Specifically, while both LTU and MolCA can be reduced to using only half the LLM with virtually no degradations, LLaVA performance seems to degrade faster after layer 8. We speculate that the major contributing factor here is down to the number of multimodal tokens. In general usage, while both LTU and MolCA compress modality information into 32 QFormer tokens, LLaVA uses all 576 vision tokens from the encoder. Consequently, this may lower redundancy due to the increase in multimodal load. Additionally, redundancy in MolCA and LTU may be more due to higher training exposures (higher ratios of data/training iterations to the number of multimodal tokens, when compared to LLaVA).
6 Conclusion & Future Work
In this work, we first analyze attention maps in multimodal models, revealing significant redundancy in their early layers. To exploit this, we introduce DeepInsert, a method validated through extensive experiments across different modalities. Our results demonstrate that DeepInsert achieves (a) substantial efficiency gains with negligible impact on performance up to a certain threshold, and (b) a favorable performance-efficiency trade-off when pushed further. The most compelling advantage of DeepInsert is its versatility: it can be seamlessly applied to off-the-shelf MLLMs without any additional hyperparameter-tuning or adaptation.
Our findings open several research directions. First, we must explore the empirical limits of DeepInsert across model sizes and dataset diversity. A natural extension is to video MLLMs, where long-form content poses substantial computational challenges. Second, rigorously characterizing why different models yield varying performance-efficiency trade-offs could advance both our mechanistic understanding of multimodal processing and practical methods for composing pretrained models. Here, we point the reader to recent concurrent work by Hartman et al. (2025), that studies this question for VLMs with careful theoretical rigor. Finally, formalizing our hypothesis that multimodal tokens function as queries (see Appendix A) may illuminate multimodal mechanisms for factual retrieval analogous to those in LLMs.
Limitations